# VideoWebArena: Evaluating Long Context Multimodal Agents with Video Understanding Web Tasks

### v2

@@ Abstract
Videos are often used to learn or extract the necessary information to complete tasks in ways different than what text or static imagery can provide. However, many existing agent benchmarks neglect long-context video understanding, instead focusing on text or static image inputs. To bridge this gap, we introduce VideoWebArena (VideoWA), a benchmark for evaluating the capabilities of long-context multimodal agents for video understanding. VideoWA consists of 2,021 web agent tasks based on manually crafted video tutorials, which total almost four hours of content. For our benchmark, we define a taxonomy of long-context video-based agent tasks with two main areas of focus: skill retention and factual retention. While skill retention tasks evaluate whether an agent can use a given human demonstration to complete a task efficiently, the factual retention task evaluates whether an agent can retrieve instruction-relevant information from a video to complete a task. We find that the best model achieves a 13.3% success rate on factual retention tasks and 45.8% on factual retention QA pairs—far below human success rates of 73.9% and 79.3%, respectively. On skill retention tasks, long-context models perform worse with tutorials than without, exhibiting a 5% performance decrease in WebArena tasks and a 10.3% decrease in VisualWebArena tasks. Our work highlights performance gaps in the agentic abilities of long-context multimodal models and provides as a testbed for the future development of long-context video agents.111 Link to code https://github.com/ljang0/videowebarena/
@@ 1 Introduction
Humans often use videos to complete daily tasks from following tutorials to retrieving information from within one or across several videos. As large foundation models are adapted into AI agents Zheng et al. (2024); Wang et al. (2023); Putta et al. (2024), these multimodal agents may also need similar capabilities to understand and process videos in order to complete various tasks—especially if they need to retrieve and process multimodal information (e.g., from the web) in order to complete said tasks. In these scenarios, the ability of models to maintain long-term memory Zhang et al. (2024b), retrieve and use said information Lewis et al. (2020), and adapt to new information continuously Wu et al. (2024c) is critical for complex, multi-step tasks that require sustained engagement over time.
@@
To address this gap, we present VideoWebArena, a novel, open-source video-based benchmark that evaluates multimodal models’ agentic ability to process, understand, and utilize long-context video inputs to accomplish various tasks. We summarize our contributions as follows.
@@
  • We evaluate popular and recently video-capable state-of-the-art LLMs (e.g., GPT-4o and Gemini 1.5 Pro) on our benchmark to provide an overview of these models’ current long-context video understanding capabilities. Our results show that these models can serve as video-capable agents, but in a limited capacity as they are still far from human levels of performance, highlighting a considerable gap in the information retrieval and agentic abilities of current state-of-the-art long-context models.
For details on the benchmark’s code and documentation, please see our reproducibility statement at the end of this paper.
@@ 2 Background
Large vision language models (VLMs) have been popular study subjects for incorporating video input and can be characterized by their learning mechanisms or overall architectural design. Popular state-of-the-art (SOTA) models like the GPT-4 (OpenAI, 2024) family of models, Claude (AI Anthropic, 2024), and Gemini (Google, 2023) are now able to handle not just text but also visual and even audio input.
Generally, similar to LLMs, VLM architectures typically revolve around two types—models with either a joint encoder-decoder architecture such as LLaVA and its variations (Liu et al., 2023) or a decoder-only architecture. Encoder-decoder VLMs tend to project different modalities through a shallow neural network or fully connected layer to link modalities. Decoder-style models typically rely on a decoder-only LLM that processes raw inputs (e.g., text tokens, image patches, etc.) such as VILA (Lin et al., 2023) and its variants like VILA2 (Fang et al., 2024) and X-VILA (Ye et al., 2024). As multimodal understanding with long-context capability becomes more important in processing more input information like video data, models like LongVILA (Xue et al., 2024) can now process a much larger number of video frames in its input for longer videos.
@@
Recent works have turned to adapting large multimodal models like VLMs into autonomous agents, given their capabilities in visual question answering and high-level reasoning. Many approaches have been further developed to improve or supplement the agentic capabilities of these large models, ranging from data collection and synthesis for specific kinds of training (Lai et al., 2024; Patel et al., 2024) or for use at inference time (Pan et al., 2024; Fu et al., 2024b; Koh et al., 2024b; Sarch et al., 2024; Wang et al., 2024). Similar works have shown how certain fine-tuning (Furuta et al., 2024) or prompting techniques (e.g., Set-of-Marks (Yang et al., 2023), (Chi et al., 2024)) can improve performance in navigating settings like web pages to accomplish tasks. Others attempt to improve how agents themselves dynamically compose/search for policies (Sodhi et al., 2024).
@@
As more works adapt LLMs and VLMs into agents, other works have focused on evaluating such efforts. These benchmarks can range across a variety of settings: from general web browsing (Yao et al., 2022; Deng et al., 2023; Zhou et al., 2024; Koh et al., 2024a), where agents are evaluated on their abilities to navigate the web and accomplish specific tasks, to mobile environments, where agents are expected to perform tasks within a mobile OS simulation like Android (Zhang et al., 2024a; Rawles et al., 2024). Other general environments attempt to emulate an OS or computing environment like MMInA (Zhang et al., 2024c), OSWorld (Xie et al., 2024), and Windows Agent Arena (Bonatti et al., 2024) where agents must navigate across multiple computer applications online and offline. Custom environments like WorkArena (Drouin et al., 2024) instead target more specific platforms like ServiceNow in constructing tasks for agent evaluation.
@@ 3.1 Summary & Overview
These domains’ websites are locally hosted since the docker images for each website are publicly available online. There is an Amazon Machine Image and instructions dedicated to hosting these websites on an EC2 instance; we refer readers to the codebase for further information. By doing this, we can make our benchmark realistic but reproducible, leveraging data and code from real and popular websites on the internet. We refer readers to WebArena (Zhou et al., 2024) and VisualWebArena (Koh et al., 2024a) for more information on each site and their setup.
@@ 3.2.1 Framework
We can define an agent’s trajectory on our tasks as a partially observable Markov decision process (POMDP) $(S,\mathcal{O},\mathcal{A},T,\mathcal{R})$ with state space $S$, observation space $\mathcal{O}$, action space $\mathcal{A}$ containing actions $a$, transition function $T:S\times\mathcal{A}\rightarrow S$, and reward function $\mathcal{R}:S\times\mathcal{A}\rightarrow\mathbb{R}$. Given current observation $o_{t}\in\mathcal{O}$, an agent generates executable action $a_{t}\in\mathcal{A}$, resulting in a new state $s_{t+1}\in S$ and a new partial observation $o_{t+1}\in\mathcal{O}$. The reward function $\mathcal{R}:S\times\mathcal{A}\rightarrow[0,1]$ returns a non-zero value at the final step if the agent state achieves the task objective and zero otherwise. We list the available reward function in Table 9.
@@ 3.3.1 Action Space
The action space for the agents in the VideoWA environment can be seen in Table 4. Agents are prompted to generate a single action from the action space at each time step. Each action is associated with Playwright Python code that automatically performs the action within the browser. The ‘[elem]’ represents the unique Set-of-Marks element that can be interacted with from the observation space provided through the environment’s JavaScript code.
@@ 3.4 Task Design
The taxonomy covers two subsets of tasks—skill retention and factual retention—inspired by real-world use cases. We illustrate the taxonomy breakdown in Figure 6. We define skill retention as the ability to learn from and use a given human demonstration to efficiently complete a task For example, using YouTube tutorials or screen recordings of expert demonstrations to learn how to perform a task is a form of skill retention. On the other hand, factual retention is the ability to retrieve information relevant to a user’s specific question/task present in a video that may not be the video’s main focus (e.g., an incidental detail). For example, one might want to buy the shoes a particular NBA player is wearing that are shown within a short duration of a much longer basketball highlights video. To complete the task, the model must extract not only information about the specific player but also their shoes, even if this information is secondary to the main content of the video.
@@
We present example tasks in Table 5 and an example of an agent on a stylized task in Figure 7. Each task has an ‘intent’, which is the objective of the task. For all of the newly created factual retention tasks, there is also an ‘intermediate_intent’, a video-based question that must be answered correctly to have the information necessary to complete the task. Each task also has an automatic evaluator function for both ‘intent’ and ‘intermediate_intent’ that returns a score of 0 or 1 based on the environment and response given by the LLM agent. Each task also has an agentic difficulty, distributed between easy, medium, and hard. The agentic difficulty for each task signifies the complexity of the action sequence needed to complete an intent successfully. For agentic difficulty, we classify a task as easy if it can be completed in 1-3 steps, medium if it can be completed in 4-9 steps, and hard if it can be completed in more than 9 steps. This classification is adopted from VisualWebArena (Koh et al., 2024a). Figure 3 provides a more detailed breakdown of task difficulty.
@@ 3.6 Factual Retention Tasks
Additionally, we provide video difficulty ratings for all intermediate intents, distributed between easy, medium, and hard. The video difficulty ratings signify the complexity of returning the correct answer for a given task’s intermediate intent. Easy tasks require returning one piece of information and can be solved with less than 3 frames, medium tasks require returning 2 to 3 things and can be solved with less than half the video, and hard tasks require returning more than 3 things and require watching more than half the video. We provide a breakdown of each task type in our benchmark in Table 5. VideoWA contains 111 unique intent templates across the 400 intents in the factual retention task set.
@@ 3.7 Task Evaluation
Each task has an ‘eval’ and intermediate ‘eval’ function. We import the automatic functionality from VisualWebArena (Koh et al., 2024a) and WebArena (Zhou et al., 2024) to evaluate our agent tasks. For the intermediate intent evaluation, we use the string-based existing functionality evaluators to assess the agent’s response to the video-based question.
@@ 6 Discussion
We present VideoWebArena, a video-based agent benchmark that tests the video capability of long-context multimodal models through an agentic lens. We define a task taxonomy of video-based agent tasks, utilizing a wide coverage of task types including skill retention and factual retention to create a comprehensive test bed for the setting of video agents. We provide 2,021 tasks that are all video-based, along with 74 manually created videos.
@@
Overall, the baseline agents do not perform well on most tasks especially relative to human performance—there is still a long ways ahead in developing intelligent agents capable of effectively utilizing and understanding video in completing regular tasks. For future work, we believe further analysis of failure modes from different video agent architectures on our benchmark may be key. We hope our environment/benchmark can help facilitate improvement and additional work on improving long-context multimodal agents.

### v3

@@ Abstract
Videos are often used as a unique source of information in learning how to perform tasks in ways different than what text or static imagery can provide. However, many existing agent benchmarks neglect long-context video understanding, instead focusing on text or static image inputs. To bridge this gap, we introduce VideoWebArena (VideoWA), a benchmark for evaluating the capabilities of long-context multimodal agents for video understanding. VideoWA consists of 2,021 web agent tasks based on manually crafted video tutorials, which total almost four hours of content. For our benchmark, we define a taxonomy of long-context video-based agent tasks with two main areas of focus: skill retention and factual retention. While skill retention tasks evaluate whether an agent can use a given human demonstration to complete a task efficiently, factual retention tasks evaluate whether an agent can retrieve instruction-relevant information from a video to complete a task. We find that the best model achieves a 13.3% success rate on factual retention tasks and 45.8% on factual retention QA pairs—far below human success rates of 73.9% and 79.3%, respectively. On skill retention tasks, long-context models perform worse with tutorials than without, exhibiting a 5% performance decrease in WebArena tasks and a 10.3% decrease in VisualWebArena tasks. Our work highlights performance gaps in the agentic abilities of long-context multimodal models when it comes to video understanding and provides as a testbed for the future development of long-context video agents.111 Link to code https://github.com/ljang0/videowebarena/
@@ 1 Introduction
Humans often use videos to complete tasks—following tutorials to retrieving information from within one, or across several, videos. As large foundation models are adapted into AI agents Zheng et al. (2024); Wang et al. (2023); Putta et al. (2024), these multimodal agents may also require similar capabilities to understand and process videos in completing various tasks—especially if they need to retrieve and process said multimodal information (e.g., from the web) to complete tasks that require new knowledge. In these scenarios, the ability of models to maintain long-term memory Zhang et al. (2024b), retrieve and use said information Lewis et al. (2020), and adapt to new information continuously Wu et al. (2024c) is critical for complex multi-step tasks.
@@
To address this gap, we present a novel, open-source video-based benchmark that evaluates multimodal models’ agentic ability to process, understand, and utilize long-context video to accomplish various tasks. In sum, our contributions are:
@@
  • We evaluate popular and recent video/image-capable state-of-the-art LLMs (e.g., GPT-4o and Gemini 1.5 Pro) to better understand their current long-context video understanding capabilities. Our results show that video/image-capable agents are still limited and far from human levels of performance, highlighting a considerable gap in the information retrieval and agentic abilities of current state-of-the-art long-context models.
The rest of our paper is organized as follows. We begin with a discussion of prior works in Section 2 before detailing the benchmark environment in Section 3. Section 4 describes our experimental setup and Section 5 lays out our benchmarking results. We conclude and discuss our findings in Section 6.
@@ 2 Background
Large vision language models (VLMs) have been popular candidates for incorporating video input understanding. Popular state-of-the-art (SOTA) models like the GPT-4 (OpenAI, 2024) family of models, Claude (AI Anthropic, 2024), and Gemini (Google, 2023) are now able to handle not just text but also visual and even audio input. Similar to LLMs, VLM architectures typically revolve around two types—models with either a joint encoder-decoder architecture such as LLaVA and its variations (Liu et al., 2023) or a decoder-only architecture. Encoder-decoder VLMs tend to project different modalities through a shallow neural network or fully connected layer to link modalities while decoder-style models typically rely on a decoder-only LLM that processes raw inputs (e.g., text tokens, image patches, etc.) such as VILA (Lin et al., 2023) and its variants like VILA2 (Fang et al., 2024) and X-VILA (Ye et al., 2024). As multimodal understanding with long-context capability becomes more important in processing increasingly large amounts of input information, like video data, models like LongVILA (Xue et al., 2024) can now process a much larger number of video frames in its input for longer videos.
@@
Recent work has turned to adapting large multimodal models like VLMs into autonomous agents, given their capabilities in visual question answering, instruction-following, and high-level reasoning. Many approaches have been developed to improve or supplement the agentic capabilities of these large models, ranging from data collection and synthesis for specific kinds of training (Lai et al., 2024; Patel et al., 2024) or for use at inference time (Pan et al., 2024; Fu et al., 2024b; Koh et al., 2024b; Sarch et al., 2024; Wang et al., 2024). Similar works have shown how certain fine-tuning (Furuta et al., 2024) or prompting (e.g., Set-of-Marks (Yang et al., 2023), (Chi et al., 2024)) can improve performance in navigating web pages to accomplish tasks. Other works have attempted to improve how agents dynamically compose/search for policies themselves (Sodhi et al., 2024).
@@
As more works adapt LLMs and VLMs into agents, there is a need to evaluate their performance. These benchmarks can range across a variety of settings: from general web browsing (Yao et al., 2022; Deng et al., 2023; Zhou et al., 2024; Koh et al., 2024a), where agents are evaluated on their abilities to navigate the web and accomplish specific tasks, to mobile environments, where agents are expected to perform tasks within a mobile OS simulation like Android (Zhang et al., 2024a; Rawles et al., 2024). Other general environments attempt to emulate an OS or computing environment like MMInA (Zhang et al., 2024c), OSWorld (Xie et al., 2024), and Windows Agent Arena (Bonatti et al., 2024) where agents must navigate across multiple computer applications online and offline. Custom environments like WorkArena (Drouin et al., 2024) instead target more specific platforms like ServiceNow in constructing tasks for agent evaluation.
@@ 3.1 Summary & Overview
These domains’ websites are locally hosted since the docker images for each website are publicly available online. There is an Amazon Machine Image and instructions dedicated to hosting these websites on an EC2 instance; we refer readers to our codebase for further information. By doing this, we can make our benchmark realistic and reproducible, leveraging data and code from real and popular websites on the internet. We refer readers to WebArena (Zhou et al., 2024) and VisualWebArena (Koh et al., 2024a) for more information on each site and its setup.
@@ 3.2.1 Framework
We can define an agent’s trajectory on our tasks as a partially observable Markov decision process (POMDP) $(S,\mathcal{O},\mathcal{A},T,\mathcal{R})$ with state space $S$, observation space $\mathcal{O}$, action space $\mathcal{A}$ containing actions $a$, transition function $T:S\times\mathcal{A}\rightarrow S$, and reward function $\mathcal{R}:S\times\mathcal{A}\rightarrow\mathbb{R}$. Given current observation $o_{t}\in\mathcal{O}$, an agent generates executable action $a_{t}\in\mathcal{A}$, resulting in a new state $s_{t+1}\in S$ and a new partial observation $o_{t+1}\in\mathcal{O}$. The reward function $\mathcal{R}:S\times\mathcal{A}\rightarrow[0,1]$ returns a non-zero value at the final step if the agent state achieves the task objective and zero otherwise. We list available rewards and evaluators in Table 9.
@@ 3.3.1 Action Space
We describe the action space of agents within the VideoWA environment in Table 4. Agents are prompted to generate a single action from the action space at each time step. Each action is associated with Playwright Python code that automatically performs the action within the browser. The ‘[elem]’ represents the unique Set-of-Marks element that can be interacted with from the observation space provided through the environment’s JavaScript code.
@@ 3.4 Task Design
The taxonomy covers two overall task types—skill retention and factual retention—inspired by real-world use cases. We illustrate the taxonomy breakdown in Figure 6. We define skill retention as the ability to learn from and use a given human demonstration to efficiently complete a task For example, using YouTube tutorials or screen recordings of expert demonstrations to learn how to perform a task is a form of skill retention. On the other hand, factual retention is the ability to retrieve information relevant to a user’s specific question/task present in a video that may not be the video’s main focus (e.g., an incidental detail). For example, one might want to buy the shoes a particular NBA player is wearing but are only shown within a short duration of a much longer basketball highlights video. To complete the task, the model must extract not only information about the specific player but also their shoes, even if this information is secondary to the main content of the video.
@@
We present example tasks in Table 5 and an example of an agent on a stylized task in Figure 7. Each task has an ‘intent’, which is the objective of the task. For all of the newly created factual retention tasks, there is also an ‘intermediate_intent’, a video-based question that must be answered correctly to have the information necessary to complete the task. Each task also has an automatic evaluator function for both ‘intent’ and ‘intermediate_intent’ that returns a score of 0 or 1 based on the environment and response given by the LLM agent. Each task also has an agentic difficulty, distributed between easy, medium, and hard. The agentic difficulty for each task signifies the complexity of the action sequence needed to complete an intent successfully. For agentic difficulty, we classify a task as easy if it can be completed in 1-3 steps, medium if it can be completed in 4-9 steps, and hard if it can be completed in more than 9 steps. This classification is adopted from VisualWebArena (Koh et al., 2024a). Figure 3 provides a detailed breakdown of task difficulty.
@@ 3.6 Factual Retention Tasks
Additionally, we provide video difficulty ratings for all intermediate intents: easy, medium, and hard. The difficulty ratings signify the complexity of returning the correct answer for a given task’s intermediate intent. Easy tasks require returning one piece of information and can be solved with less than 3 frames, medium tasks require returning/retrieving 2 to 3 pieces of information and can be solved with less than half the video, and hard tasks require returning more than 3 pieces that require watching more than half the video. We provide a breakdown Table 5. VideoWA contains 111 unique intent templates across the 400 intents in the factual retention task set.
@@ 3.7 Task Evaluation
Each task has a final and intermediate ‘eval’ function. We import the automatic functionality from VisualWebArena (Koh et al., 2024a) and WebArena (Zhou et al., 2024) to evaluate our agent tasks. For the intermediate intent evaluation, we use the string-based existing functionality evaluators to assess the agent’s response to the video-based question.
@@ 6 Discussion
We present VideoWebArena, a video-based agent benchmark that tests the video capability of long-context multimodal models through an agentic lens, consisting of 2,021 tasks that are all video-based, along with 74 manually created videos. These tasks follow a custom defined taxonomy of video-based tasks, providing a wide coverage of task types including skill retention and factual retention to create a comprehensive test bed.
@@
Overall, baseline agents do not perform well on most tasks, falling far behind human performance. It is clear that there is still a long ways ahead in developing intelligent agents capable of effectively utilizing and understanding video in completing regular tasks. As such, further analysis of failure modes from different video agent architectures on our benchmark is key. We hope our environment/benchmark can help facilitate improvement and additional work on improving long-context multimodal agents.